# o m ∥#2 ∥_#1

### v1

@@ 1 Introduction
    We describe new procedures for (a) replicable regression with random designs and (b) replicable uncentered covariance estimation, both of which may be of independent interest.
@@
    We apply these tools to develop the first replicable RL algorithms for linear MDPs, encompassing both generative models and episodic exploration.
@@ 2.1 Linear Markov decision processes
$R_{h}(s,a)=\langle\phi(s,a),\mathbf{\theta}_{h}\rangle\quad\text{and}\quad P_{h}(\cdot|s,a)=\langle\phi(s,a),\mathbf{\mu}_{h}\rangle\enspace.$
@@ 2.2 Replicability
Our results rely on a key technique by [ILPS22] that uses randomized rounding to obtain replicability in vector spaces which we extend to matrix spaces. We state a slightly adapted version of the results in Algorithm 1 and Theorem 2.1.
@@
3 Replicable ridge regression and uncentered covariance estimation
@@ 3.1 Replicable ridge regression
To effectively operate within the linear space of our MDPs, a critical first step is to develop a procedure for replicable linear regression. Prior work on bandits already provided a replicable least-squares estimator in the fixed design setting [EKKK+23] where one has access to a distribution $\nu$ of a fixed set input vectors $\mathbf{x}$. This approach is limited to scenarios in which a fixed design distribution can be obtained replicably. In addition, it assumes that the design distribution sufficiently spans the input space. Both of these conditions are not common in RL, where the distribution of visited states evolves as the agent explores its environment. Our first contribution is a novel algorithm that builds on the well-known ridge regression algorithm [HK70] to circumvent these issues. This algorithm works both in scenarios where the full space is not spanned as well as in problems with random design.
@@
Suppose Assumption 3.1 holds. Let $\varepsilon,\delta,\rho\in(0,1)$. Then for any sequence of independent distributions $D_{[t]}=\{D_{1},...,D_{t}\}$ from which we each draw $M$ independent samples totalling $N$, we have that Algorithm 2 is $\rho$-replicable. Let $\theta^{*}=\arg\min_{\theta}\mathop{\mathbb{E}}_{D_{[t]}^{M}}[(\theta^{\top}\mathbf{x}-y)^{2}+\lambda\|\theta\|_{2}^{2}]$. Then with probability at least $1-\delta$ over choice of the sample $S\sim D_{[t]}^{M}$, it holds that $\|\overline{\mathbf{w}}-\theta^{*}\|\leq\varepsilon$ as long as the number of samples drawn is $N\in\Omega\left(\frac{(B+Y)^{2}d^{3}}{\lambda^{2}\varepsilon^{2}(\rho-2\delta)^{2}}\log\left(\frac{1}{\delta}\right)\right)\enspace.$
@@
We acknowledge that the generality leads to slightly worse bounds than those achieved in the fixed design setting. This is a function of us requiring a technique that works even when the support of the data is small. If one is instead able to make assumptions about the eigenvalues of the second order moment one can recover tighter rates, close to the fixed design rates of [EKKK+23] by setting $\lambda=0$.
@@ 3.2 Replicable uncentered covariance estimation
The second tool we need is a procedure for obtaining a replicable estimate of the second order moment matrix, which we will use to identify parts of the state space that have been visited. In Algorithm 3 we provide this replicable estimation procedure. The algorithm first computes the upper-triangular part of the regular uncentered covariance, randomly rounds it, and then copies it to the lower-triangular part. This makes sure the matrix remains positive semi-definite. We guarantee replicability and closeness in Frobenius norm to the expected uncentered covariance via the following Theorem.
@@
3:Compute $\widehat{\Sigma}_{jl}=\sum_{t=0}^{T-1}\frac{1}{M}\sum_{m=0}^{M-1}\mathbf{x}_{t,m}^{j}\,\mathbf{x}_{t,m}^{l},\quad 1\leq j\leq l\leq d.$ $\triangleright$ upper-triangular half-matrix
@@
6:return $\overline{\Sigma}$
@@
Suppose Assumption 3.1 holds. Let $\varepsilon,\delta,\rho\in(0,1)$. For any sequence of independent distributions $D_{[T]}=\{D_{1},...,D_{T}\}$ from which we each draw $M$ independent samples totalling NN Algorithm  3 is $\rho$-replicable. With probability at least $1-\delta$ over the independent draw of the dataset $\mathcal{D}=\{\mathbf{x}_{t,m}\}_{(t,m)\in[T]\times[M]}$, it holds that $\|\overline{\Sigma}-\mathop{\mathbb{E}}[\mathbf{x}\mathbf{x}^{\top}]\|_{F}\leq\varepsilon$ as long as we draw $N\in\Omega\left(\frac{d^{8}T^{2}}{\varepsilon^{2}(\rho-\delta)^{2}}\log\left(\frac{d^{2}}{\delta}\right)\right)$ samples.
@@ 4.1 Replicable linear RL with generative models
As a warmup, we will consider the setting of RL with a generative model [KS98]. In this setting, one is given access to a generative model $G_{\mathcal{\mathcal{M}}}$ that given a state $s_{h}$ and an action $a_{h}$ returns a deterministic reward $R_{h}$ and a next state $s_{h+1}$ sampled from the transition probability $P_{h}(\cdot|s_{h},a_{h})$. In general results can be easily extended to stochastic rewards. In the tabular setting, it is common to simply sample every state-action pair from the environment sufficiently often until enough data for statistical concentration is available (e.g. [KS98, Kak03]). In the linear setting this is unfortunately not possible since there are possibly infinitely many state-action pairs. Instead, we need to obtain a set of representative state-action pairs that will cover the lower dimensional space of $\phi$. Such a set of vectors is often assumed given and can be represented via a set of states that gives Malahanobis distance guarantees [YW19] or via an optimal design [LS20]. Given that R-Ridge-Regression works without second order moment assumptions, we will reuse our core set as given in Definition 3.1.
@@
The accuracy proof follows the standard core set ideas [LS20]. For replicability, we know the only randomness comes from sampling the data. We start with a fixed initialization; then inductively make a replicable estimate of the prior round’s value. The full proof is provided in Appendix C.
@@ 4.2 Replicable linear RL with exploration
The previous section illustrated that it is possible to obtain replicable algorithms in the linear MDP setting. However, so far we assumed that we have access to a specific core set of state-action pairs of which we can draw next-state samples as we please. A key challenge for replicability is to deal with the noisy exploration process in RL. In the exploration setting, it is possible to obtain optimal policies that are non-replicable using LSVI-UCB [JYWJ20]. This section builds on this well-established finding and provides a replicable version of the LSVI-UCB algorithm.
The algorithm provided in Algorithm 5 proceeds in rounds. Rather than updating our policy at every episode, we collect a batch of sampled data with the current policy first to obtain replicable estimates of the data-dependent quantities required by the algorithm. The algorithm then uses R-Ridge-Regression to obtain a mapping from $\phi$ to an optimistic prediction of the value. For exploration, we add a second order moment bonus term that is computed using R-UC-Cov-Estimation. The following theorem statement highlights the guarantees we provide for the algorithm.
@@ 4.3 Limitations
The two algorithms we present are both replicable and give strong stability guarantees in a linear function approximation MDP setting. However, their sample complexity cost is larger than that of their non-replicable counterparts and linear MDPs alone are often insufficient inpractice. While recent work has shown promising results employing linear MDPs on common benchmarks [ZRYG+22], they require a meticuluous feature learning procedure that has no replicability guarantees. To take the last step towards fully practical replicability, the feature learning problem for low-rank MDPs [JKAL+17, DKWY20, AKKS20, MCKJ+24] remains an interesting open problem in terms of replicability.
@@ 5 Experimental evaluation
While some of our worst-case guarantees might seem impractical, this section shows that in practice, our algorithms require far fewer samples to work effectively. In addition, we show that even though our results are largely derived for linear MDPs with fixed feature representations, the ideas behind replicability might be valuable to study even in the non-linear deep RL setting. First, we evaluate our algorithm and its components on the well-studied CartPole environment [BSA90]. Then, we study the effects of quantized neural network Q-values in Atari environments [BNVB13].
@@ 5.1 Evaluating replicability on real datasets
To show that our algorithms do not require impractically large amounts of data, we implement a version of fitted Q-iteration [EGW05] with replicable rounding akin to our generative model algorithm. We use the offline dataset available via d3rlpy [SI22]. We use a random Fourier feature encoding for $\phi$ and use ridge regression in each round to fit the value function. We perform $5$ rounds of value backups and set the rounding bin size to $\alpha=0.2$.
We vary two components: To ensure that all policies are trained on distinct samples and to assess the amount of data needed for replicability, we sub-sample a fraction of the data for training. Then, we vary the parameter $\lambda$ to examine its impact. We evaluate the cumulative return as a measure of policy quality, and the largest fraction of identical learned weights across all runs. Figure 1 presents the results, averaged over $100$ algorithm runs.
Our results show that replicability is achieved with a fraction of the available data and that it is correlated with high returns. This suggests that when the algorithm fails to fit the value function, replication of policies becomes unlikely. While one might expect regularization to play a role, its effect appears negligible here, likely because a few weights are disproportionately large regardless of decay. Available data seems to be the driver for replicability.
5.2 Quantizing neural network predictions
@@
While we previously noted that achieving replicability in a neural network setting might be difficult without further work on feature learning, we set out to study the effects of our algorithmic elements on deep learning algorithms. To this end, we use the recent PQN algorithm [GFEP+25] to train Q-functions on Atari Breakout and MsPacman. Our theory suggests that rounding weights and using regularization can give rise to stability. The implication of rounded weights is rounded Q-values. Rather than rounding weights directly, which may lead to unforseen challenges in deep learning, we round the outputs of our neural networks onto a fixed grid. We compare a version of quantized Q-values against regular PQN as well as regularized versions of both. We measure the final return by averaging each training run’s $3$ final episode returns. Then, we take holdout offline datasets from Minari [YPBD+24] that contain expert data for the chosen tasks. On these, we compute the pairwise action agreement across seeds. We report mean and $1.96\times$ standard error over $15$ runs in Figure 2 and hyperparameters in Appendix E.
We note that regularization alone is insufficient to ensure high agreement on either task while quantization leads to increased agreement. This is in part attributable to low action gaps [Far11] in the Atari games. While the quantized policies agree, they can do so on the wrong actions as indicated by the low return on Breakout. When combining regularization and quantization, the algorithm’s return is within variance of the baseline, indicating no loss of performance. In addition, the benefits of agreement from quantization are kept providing empirical evidence for our theoretical findings.
@@ 6 Related work
Replicability     The seminal idea of algorithmic stability has a long history in learning theory and given rise to various settings such as error stability [KR99], uniform stability [BE02], or differential privacy [DMNS06], each quantifying how sensitive an algorithm’s output is to changes in its input data. Recently notions of formal reproducibility have been proposed [AJJK+22, ILPS22]. The notion we call replicability [ILPS22] was introduced to study the limits of stability. It asks that two executions of an algorithm on two different samples from the same distribution will yield the exact same outcome. Replicability is strongly related to aforementioned areas like privacy, or even generalization [BGHI+23, KKMV23]. Since its inception, replicability has been studied for clustering [EKMV+23], large-margin half spaces [KKLV+24], hypothesis testing [LY24, HIKL+24, AACN+25], geometric partitions [VDPR+24], and online settings [EKKK+23, ABB24, KIYK24]. [HM25] study the role of randomness for replicability and [KKVZ24] provide an overview over the computational landscape of replicability, relating it to other existing areas. Closely related to ours is the work on replicable tabular RL [EHKS23, KVYZ23, HLYY25].
@@ A.2 Gradient concentration and strong convexity
Because $R(\theta)$ is $2\lambda$-strongly convex, it has a unique minimizer $\tilde{\theta}$. Conditioned on the fact that
@@
$\Longleftrightarrow\quad$ $\varepsilon^{\prime}=\frac{2\lambda\varepsilon(\rho-\delta)}{d^{3/2}+\rho-2\delta}$
@@
$\max_{\mathbf{x}}|\mathbf{x}^{T}(\hat{\theta}-\theta^{*})|\leq\varepsilon.$
@@
$|\mathbf{x}^{T}(\hat{\theta}-\theta^{*})|\leq|\mathbf{x}^{T}(\hat{\theta}-\tilde{\theta})|+|\mathbf{x}^{T}(\tilde{\theta}-\theta^{*})|$
By data assumption and our Theorem 3.2, we can have that $|\mathbf{x}^{T}(\hat{\theta}-\tilde{\theta})|\leq\|x\|\|(\hat{\theta}-\tilde{\theta})\|\leq\cfrac{\varepsilon}{2}$. It remains to prove that $|\mathbf{x}^{T}(\tilde{\theta}-\theta^{*})|$ is small. We can use the core-set assumption to rewrite this term as follows
@@
$=\sum_{i\in[t]}\sum_{m\in[M]}\phi(s^{i}_{m,h},a^{i}_{m,h})(R_{m,h}^{i}+\hat{V}_{h+1}^{t}(s^{i}_{m,h+1}))-\sum_{i\in[t]}\sum_{m\in[M]}\phi(s^{i}_{m,h},a^{i}_{m,h})(R^{i}_{m,h}+\mathop{\mathbb{E}}_{s^{\prime}\sim P^{i}_{m,h}}[V_{h+1}^{\pi}(s^{\prime})])-\lambda\mathbf{w}_{h}^{\pi}$
@@
$=\sum_{i\in[t]}\sum_{m\in[M]}\phi(s^{i}_{m,h},a^{i}_{m,h})(\hat{V}^{t}_{h+1}(s^{i}_{m,h+1})+\mathop{\mathbb{E}}_{s^{\prime}\sim P^{i}_{m,h}}[\hat{V}^{t}_{h+1}(s^{\prime})-\hat{V}^{t}_{h+1}(s^{\prime})-V_{h+1}^{\pi}(s^{\prime})])-\lambda\mathbf{w}_{h}^{\pi}$
@@
$\leq\|\phi(s,a)\|_{(\Lambda_{h}^{t})^{-1}}\|\sum_{i\in[k]}\sum_{m\in[M]}\phi(s^{i}_{m,h},a^{i}_{m,h})(\hat{V}^{t}_{h+1}(s^{i}_{m,h+1})-\mathop{\mathbb{E}}_{s^{\prime}\sim P^{i}_{m,h}}[\hat{V}^{t}_{h+1}(s^{\prime})])\|_{(\Lambda_{h}^{t})^{-1}},$
so it suffices to bound $\|\sum_{i\in[k]}\sum_{m\in[M]}\phi(s^{t}_{m,h},a^{t}_{m,h})(\hat{V}^{t}_{h+1}(s^{t}_{m,h+1})-\mathop{\mathbb{E}}_{s^{\prime}\sim P^{t}_{m,h}}[\hat{V}^{t}_{h+1}(s^{\prime})])\|_{(\Lambda_{h}^{t})^{-1}}$.
@@
$=\langle\phi(s,a),\int\hat{V}^{t}_{h+1}(s^{\prime})-V^{\pi}_{h+1}(s^{\prime})d\mu_{h}(s^{\prime})\rangle-\langle\phi(s,a),\lambda(\Lambda_{h}^{t})^{-1}\int\hat{V}^{t}_{h+1}(s^{\prime})-V^{\pi}_{h+1}(s^{\prime})d\mu_{h}(s^{\prime})\rangle$
@@
$\geq Q^{\pi}_{h}(s,a)-(H-h+1)O(\tfrac{\varepsilon}{H})$ by inductive assumption
@@
$\tfrac{1}{M}\sum_{t\in[T]}\sum_{h\in[H]}\sum_{m\in[M]}\|\phi(s^{t+1}_{m,h},a^{t+1}_{m,h})\|^{2}_{(\hat{G}^{t}_{h})^{-1}}$ $\leq(1+1/\lambda)\sum_{t\in[T]}\sum_{h\in[H]}\ln(1+\tfrac{1}{M}\sum_{m\in[M]}\|\phi(s^{t+1}_{m,h},a^{t+1}_{m,h})\|^{2}_{(\hat{G}^{t}_{h})^{-1}})$
@@
$\quad\quad\quad+\sum_{t\in[T]}\sum_{m\in[M]}\sum_{h\in[H]}\beta\|\phi(s^{t}_{m,h},a^{t}_{m,h})\|_{(\Lambda^{t}_{h})^{-1}}+\beta\|\phi(s^{t}_{m,h},a^{t}_{m,h})\|_{(\bar{\Lambda}^{t}_{h})^{-1}}+O(\Delta_{w})$

### v2

# Replicable Reinforcement Learning with Linear Function Approximation

@@ 1 Introduction
    We describe new procedures with first guarantees for (a) replicable regression with random designs and (b) replicable uncentered covariance estimation, both of which may be of independent interest.
@@
    We apply these tools to develop the first replicable RL algorithms for linear MDPs, encompassing both the generative model and episodic exploration setting.
@@ 2.1 Linear Markov decision processes
$R_{h}(s,a)=\langle\phi(s,a),\mathbf{\theta}_{h}\rangle\quad\text{and}\quad P_{h}(\cdot|s,a)=\langle\phi(s,a),\mathbf{\mu}_{h}\rangle\kern 5.0pt.$
@@ 2.2 Replicability
Our results rely on a key technique by [ILPS22] that uses randomized rounding to obtain replicability in vector spaces which we extend to matrix spaces. We state a slightly adapted version of their results in Algorithm 1 and Theorem 2.1.
@@
3 Replicable Tools for Linear Spaces
@@ 3.1 Replicable ridge regression
To effectively operate within the linear MDP space, a first step is to develop a procedure for replicable linear regression. Prior work on bandits provided a replicable least-squares estimator in the fixed design setting [EKKK+23] where one has access to a distribution $\nu$ of a fixed set input vectors $\mathbf{x}$. This approach is limited to scenarios in which a fixed design distribution can be obtained replicably. In addition, it assumes that the design distribution sufficiently spans the input space. Both of these conditions are not common in RL, where the distribution of visited states evolves as the agent explores its environment. Our first contribution is a novel algorithm that builds on the well-known ridge regression algorithm [HK70] to circumvent these issues. This algorithm works both in scenarios where the full space is not spanned as well as in problems with random design.
@@
Suppose Assumption 3.1 holds. Let $\varepsilon,\delta,\rho\in(0,1)$. Then for any sequence of independent distributions $D_{[t]}=\{D_{1},...,D_{t}\}$ from which we each draw $M$ independent samples totalling $N$, we have that Algorithm 2 is $\rho$-replicable. Let $\theta^{*}=\arg\min_{\theta}\mathop{\mathbb{E}}_{D_{[t]}^{M}}[(\theta^{\top}\mathbf{x}-y)^{2}+\lambda\|\theta\|_{2}^{2}]$. Then with probability at least $1-\delta$ over choice of the sample $S\sim D_{[t]}^{M}$, it holds that $\|\overline{\mathbf{w}}-\theta^{*}\|\leq\varepsilon$ as long as the number of samples drawn is $N\in\Omega\left(\frac{(B+Y)^{2}d^{3}}{\lambda^{2}\varepsilon^{2}(\rho-2\delta)^{2}}\log\left(\frac{1}{\delta}\right)\right)\kern 5.0pt.$
@@
We acknowledge that the generality leads to slightly worse bounds than those achieved in the fixed design setting. This is because we require a technique that works even when the data support is small. If one is instead able to make assumptions about the eigenvalues of the second order moment one can recover tighter rates, close to the fixed design rates of [EKKK+23] by setting $\lambda=0$.
@@ 3.2 Replicable uncentered covariance estimation
The second tool we need is a procedure for obtaining a replicable estimate of the second order moment matrix, which we will use to identify parts of the state space that have been visited. In Algorithm 3 we provide this replicable estimation procedure. Two features of a covariance matrix are that it is symmetric and positive semidefinite (PSD). Simply applying element-wise randomized rounding to the regular covariance matrix might lead to an output matrix that is neither symmetric nor PSD. Thus, our algorithm first computes the upper-triangular part of the regular uncentered covariance, randomly rounds it, and then symmetrizes explicitly. Finally, we project the matrix back onto the original cone by clipping its eigenvalues to ensure the algorithm’s output is PSD. We guarantee replicability and closeness in Frobenius norm to the expected uncentered covariance via the following Theorem.
@@
3:Compute $\widehat{\Sigma}_{jl}=\sum_{t=0}^{T-1}\frac{1}{M}\sum_{m=0}^{M-1}\mathbf{x}_{t,m}^{j}\,\mathbf{x}_{t,m}^{l},\quad 1\leq j\leq l\leq d.$
@@
6:Return $\Pi_{\mathrm{PSD}}(\overline{\Sigma})$ $\triangleright$ $\Pi_{\mathrm{PSD}}(A)=U\,\operatorname{diag}(\max(\zeta,0))\,U^{\top}$ for $A=U\,\operatorname{diag}(\zeta)\,U^{\top}$
@@
Suppose Assumption 3.1 holds. Let $\varepsilon,\delta,\rho\in(0,1)$. For any sequence of independent distributions $D_{[T]}=\{D_{1},...,D_{T}\}$ from which we each draw $M$ independent samples totalling NN. Algorithm  3 is $\rho$-replicable. With probability at least $1-\delta$ over the independent draw of the dataset $\mathcal{D}=\{\mathbf{x}_{t,m}\}_{(t,m)\in[T]\times[M]}$, it holds that $\|\Pi_{\mathrm{PSD}}(\overline{\Sigma})-\mathop{\mathbb{E}}[\mathbf{x}\mathbf{x}^{\top}]\|_{F}\leq\varepsilon$ as long as we draw $N\in\Omega\left(\frac{d^{8}T^{2}}{\varepsilon^{2}(\rho-\delta)^{2}}\log\left(\frac{d^{2}}{\delta}\right)\right)$ samples.
@@ 4.1 Replicable linear RL with generative models
As a warmup, we will consider the setting of RL with a generative model [KS98]. In this setting, one is given access to a generative model $G_{\mathcal{\mathcal{M}}}$ that given a state $s_{h}$ and an action $a_{h}$ returns a deterministic reward $R_{h}$ and a next state $s_{h+1}$ sampled from the transition probability $P_{h}(\cdot|s_{h},a_{h})$. In the tabular setting, it is common to simply sample every state-action pair from the environment sufficiently often until enough data for statistical concentration is available (e.g. [KS98, Kak03]). In the linear setting this is unfortunately not possible since there are possibly infinitely many state-action pairs. Instead, we need to obtain a set of representative state-action pairs that will cover the lower dimensional space of $\phi$. Such a set of vectors is often assumed given and can be represented via a set of states that gives Malahanobis distance guarantees [YW19] or via an optimal design [LS20]. Given that R-Ridge-Regression works without second order moment assumptions, we will reuse our core set as given in Definition 3.1.
@@
The accuracy proof follows the standard core set ideas [LS20]. For replicability, we know the only randomness comes from sampling. We start with a fixed initialization; then inductively make a replicable estimate of the prior round’s value. Find the full proof in Appendix C.
@@ 4.2 Replicable linear RL with exploration
The previous section illustrated that it is possible to obtain replicable algorithms in the linear MDP setting. However, so far we assumed that we have access to a specific core set of state-action pairs of which we can draw next-state samples as we please. A key challenge for replicability is to deal with the noisy exploration process in RL. In the exploration setting, it is possible to obtain optimal policies that are non-replicable using LSVI-UCB [JYWJ20]. This section builds on this well-established finding and Algorithm 5 provides a replicable version of LSVI-UCB called $\mathtt{R\text{-}LSVI\text{-}UCB}$.
$\mathtt{R\text{-}LSVI\text{-}UCB}$ proceeds in rounds. Rather than updating the policy at every episode, it collects a batch of sampled data with the current policy to obtain replicable estimates of the required data-dependent quantities. The algorithm then uses R-Ridge-Regression to obtain a mapping from $\phi$ to a prediction of $Q$. For exploration, we add a second order moment bonus term computed via R-UC-Cov-Estimation. The following statement outlines the guarantees of $\mathtt{R\text{-}LSVI\text{-}UCB}$.
@@ 4.3 Limitations
The two algorithms we present both give strong stability guarantees in a linear function approximation MDP setting. However, their sample complexity cost is larger than that of their non-replicable counterparts and linear MDPs alone are often insufficient inpractice. While recent work has shown promising results employing linear MDPs on common benchmarks [ZRYG+22], they require a meticuluous feature learning procedure that has no replicability guarantees. Towards fully practical replicability, the feature learning problem for low-rank MDPs [JKAL+17, DKWY20, AKKS20, MCKJ+24] remains an interesting open question.
@@ 5 Experimental evaluation
While some of our worst-case guarantees might seem impractical, this section shows that in practice, our algorithms need far fewer samples to work effectively. We also show that even though our results are largely derived for linear MDPs with fixed feature representations, the ideas behind replicability might be valuable to study even in the non-linear deep RL setting. First, we evaluate our algorithm and its components on the well-studied CartPole environment [BSA90]. Then, we study the effects of quantized neural network Q-values in Atari environments [BNVB13].
@@ 5.1 Evaluating replicability on real datasets
To show that our algorithms do not require impractically large amounts of data, we implement a version of fitted Q-iteration [EGW05] with replicable rounding akin to our generative model algorithm. We use the offline CartPole dataset available via d3rlpy [SI22] and a random Fourier feature encoding for $\phi$. Over $5$ rounds, we use ridge regression to fit the value function. The the rounding bin size is $\alpha=0.2$.
We vary two components: To ensure that all policies are trained on distinct samples and to assess the amount of data needed for replicability, we sub-sample a fraction of the data for training. Then, we vary $\lambda$ to examine its impact. We evaluate the cumulative return as a measure of policy quality, and the largest fraction of identical learned weights across all runs. Figure 1 presents the results, averaged over $100$ algorithm runs.
Our results show that replicability is achieved with a fraction of the available data and is correlated with high returns. This suggests that when the algorithm fails to fit the values, replication of policies becomes unlikely. While we expect regularization to play a role, its effect appears negligible here, likely because a few weights are disproportionately large. Available data seems to be the driver for replicability.
5.2 Quantizing neural Q-values
@@
While we previously noted that achieving replicability in a neural network setting might be difficult without further work on feature learning, we set out to study the effects of our algorithmic elements on deep learning algorithms. We use the recent PQN algorithm [GFEP+25] to train Q-functions on Atari Breakout and MsPacman. Our theory suggests that rounding weights and using regularization can give rise to stability. The implication of rounded weights is rounded Q-values. Rather than rounding weights directly, which may lead to unforseen challenges in deep learning, we round the outputs of our neural networks onto a fixed grid. We compare a version of quantized Q-values against regular PQN as well as regularized versions of both. We measure the final return by averaging each training run’s $3$ final episode returns. Then, we take holdout expert datasets from Minari [YPBD+24] for the chosen tasks. On these, we compute the pairwise action agreement across seeds. We report mean and $1.96\times$ standard error over $15$ runs in Figure 2 and hyperparameters in Appendix E.
Tegularization alone is insufficient to ensure high agreement on either task while quantization leads to increased agreement. This is in part attributable to low action gaps [Far11] in the Atari games. While the quantized policies agree, they can do so on the wrong actions as indicated by the low return on Breakout. When combining regularization and quantization, the algorithm’s return is within variance of the baseline, indicating no loss of performance. In addition, the benefits of agreement from quantization are kept providing empirical evidence for our theoretical findings.
@@ 6 Related work
Replicability     The seminal idea of algorithmic stability has a long history in learning theory and given rise to various settings such as error stability [KR99], uniform stability [BE02], or differential privacy [DMNS06], each quantifying how sensitive an algorithm’s output is to changes in its input data. Recently notions of formal reproducibility have been proposed [AJJK+22, ILPS22]. The notion we call replicability [ILPS22] was introduced to study the limits of stability. It asks that two executions of an algorithm on two different samples from the same distribution will yield the exact same outcome. Replicability is strongly related to aforementioned areas like privacy, or even generalization [BGHI+23, KKMV23]. Since its inception, replicability has been studied for clustering [EKMV+23], large-margin half spaces [KKLV+24], hypothesis testing [LY24, HIKL+24, AACN+25], geometric partitions [VDPR+24], and online settings [EKKK+23, ABB24, KIYK24]. [HM25] study the role of randomness for replicability and [KKVZ24] provide an overview of the computational landscape of replicability. Closely related to ours is the work on replicable tabular RL [EHKS23, KVYZ23, HLYY25].
@@ A.2 Gradient concentration and strong convexity
Because $R(\theta)$ is $2\lambda$‐strongly convex, it has a unique minimizer $\tilde{\theta}$. Conditioned on the fact that
@@
$\Longleftrightarrow\quad$ $\varepsilon^{\prime}=\frac{2\lambda\varepsilon(\rho-2\delta)}{d^{3/2}+\rho-2\delta}$
@@
$\max_{\mathbf{x}}|\mathbf{x}^{T}(\overline{\mathbf{w}}-\theta^{*})|\leq\varepsilon.$
@@
$|\mathbf{x}^{T}(\hat{\theta}-\theta^{*})|\leq|\mathbf{x}^{T}(\overline{\mathbf{w}}-\tilde{\theta})|+|\mathbf{x}^{T}(\tilde{\theta}-\theta^{*})|$
By data assumption and our Theorem 3.2, we can have that $|\mathbf{x}^{T}(\overline{\mathbf{w}}-\tilde{\theta})|\leq\|x\|\|(\overline{\mathbf{w}}-\tilde{\theta})\|\leq\cfrac{\varepsilon}{2}$. It remains to prove that $|\mathbf{x}^{T}(\tilde{\theta}-\theta^{*})|$ is small. We can use the core-set assumption to rewrite this term as follows
@@
$\|\Pi_{\mathrm{PSD}}(\overline{\Sigma})-\Sigma\|_{F}\leq\varepsilon$
Note that the eigenvalue clipping is simply the metric projection onto the PSD cone $\mathbb{S}_{+}^{d}$ in the Hilbert space $(\mathbb{S}^{d},\langle\cdot,\cdot\rangle_{F})$ that the original covariance lies in. Metric projections onto closed convex sets in Hilbert spaces are nonexpansive (i.e., $1$-Lipschitz; see, e.g., [BC17]). In addition, since $\Sigma\in\mathbb{S}_{+}^{d}$, our clipping operator would not change the true covariance matrix at all if applied and we have $\Pi_{\mathrm{PSD}}(\Sigma)=\Sigma$. As a result, we have that
$\|\Pi_{\mathrm{PSD}}(\overline{\Sigma})-\Sigma\|_{F}=\|\Pi_{\mathrm{PSD}}(\overline{\Sigma})-\Pi_{\mathrm{PSD}}(\Sigma)\|_{F}\leq\|\overline{\Sigma}-\Sigma\|_{F}$
Thus, it is sufficient to show that the rounded matrix before projection does not incur large error. We do this by decomposing as follows.
@@
$=\sum_{i\in[t]}\sum_{m\in[M]}\phi(s^{i}_{m,h},a^{i}_{m,h})(R_{m,h}^{i}+\hat{V}_{h+1}^{t}(s^{i}_{m,h+1}))$
$\quad\quad\quad-\sum_{i\in[t]}\sum_{m\in[M]}\phi(s^{i}_{m,h},a^{i}_{m,h})(R^{i}_{m,h}+\mathop{\mathbb{E}}_{s^{\prime}\sim P^{i}_{m,h}}[V_{h+1}^{\pi}(s^{\prime})])-\lambda\mathbf{w}_{h}^{\pi}$
@@
$=\sum_{i\in[t]}\sum_{m\in[M]}\phi(s^{i}_{m,h},a^{i}_{m,h})(\hat{V}^{t}_{h+1}(s^{i}_{m,h+1})$
$\quad\quad\quad+\mathop{\mathbb{E}}_{s^{\prime}\sim P^{i}_{m,h}}[\hat{V}^{t}_{h+1}(s^{\prime})-\hat{V}^{t}_{h+1}(s^{\prime})-V_{h+1}^{\pi}(s^{\prime})])-\lambda\mathbf{w}_{h}^{\pi}$
@@
$\leq\|\phi(s,a)\|_{(\Lambda_{h}^{t})^{-1}}\|\sum_{i\in[k]}\sum_{m\in[M]}\phi(s^{i}_{m,h},a^{i}_{m,h})(\hat{V}^{t}_{h+1}(s^{i}_{m,h+1})$
$\quad\quad\quad-\mathop{\mathbb{E}}_{s^{\prime}\sim P^{i}_{m,h}}[\hat{V}^{t}_{h+1}(s^{\prime})])\|_{(\Lambda_{h}^{t})^{-1}},$
so it suffices to bound
$\|\sum_{i\in[k]}\sum_{m\in[M]}\phi(s^{t}_{m,h},a^{t}_{m,h})(\hat{V}^{t}_{h+1}(s^{t}_{m,h+1})-\mathop{\mathbb{E}}_{s^{\prime}\sim P^{t}_{m,h}}[\hat{V}^{t}_{h+1}(s^{\prime})])\|_{(\Lambda_{h}^{t})^{-1}}$.
@@
$=\langle\phi(s,a),\int\hat{V}^{t}_{h+1}(s^{\prime})-V^{\pi}_{h+1}(s^{\prime})d\mu_{h}(s^{\prime})\rangle$
$\quad\quad\quad-\langle\phi(s,a),\lambda(\Lambda_{h}^{t})^{-1}\int\hat{V}^{t}_{h+1}(s^{\prime})-V^{\pi}_{h+1}(s^{\prime})d\mu_{h}(s^{\prime})\rangle$
@@
$\geq Q^{\pi}_{h}(s,a)-(H-h+1)O(\tfrac{\varepsilon}{H})$ by induction
@@
$\tfrac{1}{M}\sum_{t\in[T]}\sum_{h\in[H]}\sum_{m\in[M]}$ $\|\phi(s^{t+1}_{m,h},a^{t+1}_{m,h})\|^{2}_{(\hat{G}^{t}_{h})^{-1}}$
$\leq(1+1/\lambda)\sum_{t\in[T]}\sum_{h\in[H]}\ln(1+\tfrac{1}{M}\sum_{m\in[M]}\|\phi(s^{t+1}_{m,h},a^{t+1}_{m,h})\|^{2}_{(\hat{G}^{t}_{h})^{-1}})$
@@
$\quad\quad\quad+\sum_{t\in[T]}\sum_{m\in[M]}\sum_{h\in[H]}\beta\|\phi(s^{t}_{m,h},a^{t}_{m,h})\|_{(\Lambda^{t}_{h})^{-1}}$
$\quad\quad\quad+\beta\|\phi(s^{t}_{m,h},a^{t}_{m,h})\|_{(\bar{\Lambda}^{t}_{h})^{-1}}+O(\Delta_{w})$